Amount of taxes payed on $125000 is federal 50% and state 30%. The federal tax is computed after state tax paid. The state tax is computed after federal tax is paid. How much is the federal and state taxes?
Solution: Let y be the total income
Therefore, State Tax= y x 30%
=30y/100
=3y/10
Income after State Tax = y - 3y/10
= 7y/10
Now, Federal Tax= 7y/10 x 50%
= 7y/20
so total tax= 3y/10 + 7y/20
Now this total tax must be equal to 125000
Therefore,
3y/10 + 7y/20 = 125000
(6y + 7y)/20 = 125000
13y/20 = 125000
so, y = (125000 x 20)/13
y= $192307.69
State Tax= $192307.69 x 30%
=$ 57692.31
Income after state tax= $192307.69 - $ 57692.31
= $ 134615.38
So, Federal Tax = $134615.38 x 50%
= $ 67307.69

Answer by Adrish(44)

Amount of taxes payed on $125000 is federal 50% and state 30%. The federal tax is computed after state tax paid. The state tax is computed after federal tax is paid. How much is the federal and state taxes?
Solution: Let y be the total income
Therefore, State Tax= y x 30%
=30y/100
=3y/10
Income after State Tax = y - 3y/10
= 7y/10
Now, Federal Tax= 7y/10 x 50%
= 7y/20
so total tax= 3y/10 + 7y/20
Now this total tax must be equal to 125000
Therefore,
3y/10 + 7y/20 = 125000
(6y + 7y)/20 = 125000
13y/20 = 125000
so, y = (125000 x 20)/13
y= $192307.69
State Tax= $192307.69 x 30%
=$ 57692.31
Income after state tax= $192307.69 - $ 57692.31
= $ 134615.38
So, Federal Tax = $134615.38 x 50%
= $ 67307.69